# The statistical thermodynamics of generative diffusion models: Phase transitions, Symmetry breaking and critical instability

### v2

@@ Abstract
Generative diffusion models have achieved spectacular performance in many areas of generative modeling. While the fundamental ideas behind these models come from non-equilibrium physics, variational inference and stochastic calculus, in this paper we show that many aspects of these models can be understood using the tools of equilibrium statistical mechanics. Using this reformulation, we show that generative diffusion models undergo second-order phase transitions corresponding to symmetry breaking phenomena. We show that these phase-transitions are always in a mean-field universality class, as they are the result of a self-consistency condition in the generative dynamics. We argue that the critical instability that arises from the phase transitions lies at the heart of their generative capabilities, which is characterized by a set of mean field critical exponents. Furthermore, using the statistical physics of disordered systems, we show that memorization can be understood as a form of critical condensation corresponding to a disordered phase transition. Finally, we show that the dynamic equation of the generative process can be interpreted as a stochastic adiabatic transformation that minimizes the free energy while keeping the system in thermal equilibrium.
@@ 1 Introduction
Generative modeling is a sub-field of machine learning concerned with the automatic generation of structured data such as images, videos and written language (Bond-Taylor et al.,, 2021). Generative diffusion models (Sohl-Dickstein et al.,, 2015), also known as score-based models, form a class of deep generative models that have demonstrated high performance in image (Ho et al.,, 2020; Song et al., 2021b, ), sound (Chen et al.,, 2020; Kong et al.,, 2020; Liu et al.,, 2023) and video generation (Ho et al.,, 2022; Singer et al.,, 2022). Diffusion models were first introduced in analogy with the physics of non-equilibrium statistical physics. The fundamental idea is to formalize generation as the probabilistic inverse of a forward stochastic process that gradually turns the target distribution $\phi(\boldsymbol{y})$ into a simple base distribution such as Gaussian white noise (Sohl-Dickstein et al.,, 2015; Song et al., 2021a, ). Recently, several works suggested that many of the dynamical properties of generative diffusion models can be understood using concepts such as spontaneous symmetry breaking (Raya and Ambrogioni,, 2023; Biroli and Mézard,, 2023; Biroli et al.,, 2024), and phase transitions (Biroli et al.,, 2024; Sclocchi et al.,, 2024). These theoretical and experimental results suggest a deep connection between generative diffusion and equilibrium statistical physics. However, these works do not directly use an equilibrium formulation, which makes it difficult to directly translate results and techniques from one field to the other. In this paper using theoretical arguments, we will show that generative diffusion models can be understood and analyzed using the tools of equilibrium statistical physics. Specifically, we define a family of Boltzmann distributions over the noise-free states, which are interpreted as (unobservable) microstates during the diffusion process. Using this reformulation, we show that generative diffusion models can undergo second-order phase transitions of the mean-field type, corresponding the the generative spontaneous symmetry breaking fist discussed in (Raya and Ambrogioni,, 2023) and further studied in Biroli et al., (2024); Li and Chen, (2024) and in Sclocchi et al., (2024). In our analysis, the time dependent variance plays the role of a temperature parameter. We obtain a self-consistent equation of state for the system, which corresponds to the fixed-point equation of the generative dynamics. We also show that the generative dynamics minimizes a regularized Helmholtz free energy, which connects diffusion models to other energy-based models in the machine learning and theoretical neuroscience literature (Friston,, 2010). Finally, similarly to what was done in Lucibello and Mézard, (2024) for associative networks, we use the statistical physics of disordered systems to characterize the transition between generalization and memorization as a disordered phase transition. To validate the theoretical predictions, we present an overview of the recent theoretical and experimental literature discussing symmetry breaking phenomena and phase transitions in trained diffusion models.
@@ 2 Preliminaries on generative diffusion models
where the expectation is taken with respect to the target distribution $\phi(\boldsymbol{y})$. A generative model can then be obtained by "inverting" Eq. 1. The inverse equation equation is
@@ 2.1 Training diffusion models as denoising autoencoders
$\nabla\log{p_{t}(\boldsymbol{x})}\approx-(t\sigma^{2})^{-1}f^{*}(\boldsymbol{x}(t),t)~{}.$ (5)
@@ 3 Diffusion models as systems in equilibrium
As defined above by the forward and inverse Langevin equation, generative diffusion models are non-equilibrium systems. In this section, we shall show that they can be reformulated using the language of equilibrium statistical mechanics, which will allow us to analyze their thermodynamics properties. As we shall see, the generative process can then be derived as an adiabatic transformation that keeps the thermodynamic system in equilibrium.
The starting point for any model in statistical mechanics is the definition of the relevant microstates. In physics, the microstate is considered an unobservable quantity. If we can observe a noise-corrupted data $\boldsymbol{x}^{*}(t)$, the obvious unobservable quantity of interest is the noise-free initial state $\boldsymbol{x}^{*}(0)=\boldsymbol{y}^{*}$. The next step is to define a Hamiltonian function on the set of microstates and all values of the time $t$. We can do this by considering the conditional probability of the data $\boldsymbol{y}$ given a noisy state $\boldsymbol{x}_{t}$:
$p(\boldsymbol{y}\mid\boldsymbol{x},t)=\frac{1}{p(\boldsymbol{x},t)}e^{-\frac{{\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert}_{2}^{2}}{2t\sigma^{2}}-\log{\phi(\boldsymbol{y})}}=\frac{1}{Z(\boldsymbol{x},t)}e^{-\mathcal{H}(\boldsymbol{y};\boldsymbol{x},t)}~{},$ (6)
@@
The statistical properties of this ensemble are crucial for diffusion models as the determine the score function, which can now be expressed as a Boltzmann average:
$\nabla\log{p_{t}(\boldsymbol{x})}=\beta(\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}-\boldsymbol{x})~{},$ (9)
where
$\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}=\int\boldsymbol{y}~{}\frac{e^{-\mathcal{H}(\boldsymbol{y};\boldsymbol{x},t)}}{Z(\boldsymbol{x},t)}\text{d}\boldsymbol{y}~{}.$ (10)
Intuitively, this equation tells us that the score vector directs the system towards the posterior average $\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}$. This leads to radically different kinds of behaviors depending on the concentration of the Boltzmann weights. As we shall see, studying the thermodynamics determined by these weights will allow us to understand several important qualitative and quantitative features of the generative dynamics. For example, as we will see in later sections, after a condensation phase transition the score will only depend on a small number of data-points, which can be detected by studying the concentration of the weights.
The thermodynamic system defined by Eq. 6 does not have a true temperature parameter. However, the quantity $\sigma t$ plays a very similar role to temperature in classical statistical mechanics. Moreover, in the Hamiltonian given by Eq .7, the dynamic variable $\boldsymbol{x}$ is analogous to then external field term in magnetic systems, which can bias the distribution of microstates towards the patterns ’aligned’ in its direction. We can imagine $\boldsymbol{x}$ as being a "slower" thermodynamic variable that interacts (adiabatically) with the statistics of the microstates.
@@ 4 Free energy, magnetization and order parameters
$\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}=-\nabla\mathcal{F}(\boldsymbol{x},t)~{}.$ (18)
This formula suggests an analogy between diffusion models and magnetic systems in statistical physics. The noisy state $\boldsymbol{x}$ can be interpreted as an external magnetic field, which induces the state of "magnetization" $\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}$. In this analogy, a diffusion model is magnetized when its distribution is biased towards a sub-set of the possible microstates.
In physics, the ’external field’ variable $\boldsymbol{x}$ is usually assumed to be controlled by the experimenter. On the other hand, in generative diffusion models $\boldsymbol{x}$ is a dynamic variable that, under the reversed dynamics, is itself attracted towards $\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}$ by the drift term:
$\nabla\log{p_{t}(\boldsymbol{x})}\propto\left(\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}-\boldsymbol{x}\right)$ (19)
In other words, if we ignore the effect of the dispersion term, the state of the system is driven towards self-consistent points where $\boldsymbol{x}$ is equal to $\langle\boldsymbol{y}\rangle_{\boldsymbol{x}}$. It is therefore interesting to study the self-consistency equation
@@
Readers familiar with statistical physics will recognize that Eq. 20 is formally identical to the self-consistency conditions used in the mean field approximation, where the external field term in one location is assumed to be determined by the magnetization of all other locations. However, it is important to note that in the case of a diffusion model, this self-consistent coupling is not approximate, as it is a natural consequence of the dynamics. Nevertheless, the formal analogy implies that the thermodynamics of generative diffusion models is formally identical to the thermodynamics of mean field models.
@@ 4.1 The susceptibility matrix
$\chi_{ij}(t)=\frac{\partial m_{i}}{\partial h_{j}}\bigg{\rvert}_{\boldsymbol{h}=0}~{},$ (22)
which tells us how sensitive the expected value is to changes in the noisy-state $\boldsymbol{x}(t)$. The susceptibility matrix has a great value in interpreting the dynamics of the generative denoising process, as it informs us on how random fluctuations in each component of the state $\boldsymbol{x}(t)$ are propagated to the other components. For example, in the context of image generation, a random fluctuation of "green" at the bottom of an image will propagate to the rest, originating the image of a forest.
4.2 The regularized free energy
where $\tilde{\mathcal{F}}$ is the free-energy plus a $L_{2}$ regularization term:
~$\tilde{\mathcal{F}}(\boldsymbol{x},t)=\mathcal{F}(\boldsymbol{x},t)+\frac{1}{2}{\left\lVert\boldsymbol{x}\right\rVert}_{2}^{2}$ (24)
This can be seen as a form of adiabatic approximation, where the dynamics of the ’slow’ variable $\boldsymbol{x}$ is obtained by assuming that the system is maintained in thermal equilibrium along the diffusion trajectory. The symmetry breaking can now be detected as a change of shape in the regularized free energy, which transitions from a convex shape with a single global minimum to a more complex shape with potentially several meta-stable points (see Fig. 3). The reformulation of the dynamics in term of the gradient of a free energy allows us to interpret generative diffusion models as a kind of energy-based machine learning models (LeCun et al.,, 2006), as discussed in (Hoover et al.,, 2023) and (Ambrogioni,, 2023). The main difference is that the (free) energy is not learned directly but it is instead implicit in the learned score function. The connection suggests potential connection with the free energy principle in theoretical neuroscience, which is used to characterize the stochastic dynamics of biological neural systems (Friston,, 2010).
@@ 4.2 The regularized free energy
5 Connecting the thermodynamics of diffusion models and physical systems: The diffused Ising model
@@ 6 Phase transitions and symmetry breaking
A spontaneous symmetry breaking happens when this trivial solution for the order parameter branches into multiple solutions at some critical time of $t_{c}$. This corresponds to the onset of multi-modality in the regularized free energy ~$\tilde{\mathcal{F}}(\boldsymbol{x},t)$, as visualized in Fig. 3 for a "four deltas" model. In thermodynamic systems, the symmetry breaking corresponds to a (second order) phase transition, which in this case can be detected by the divergence of several state variables around the critical point.

### v4

@@ Abstract
Generative diffusion models have achieved spectacular performance in many areas of machine learning and generative modeling. While the fundamental ideas behind these models come from non-equilibrium physics, variational inference and stochastic calculus, in this paper we show that many aspects of these models can be understood using the tools of equilibrium statistical mechanics. Using this reformulation, we show that generative diffusion models undergo second-order phase transitions corresponding to symmetry breaking phenomena. We show that these phase-transitions are always in a mean-field universality class, as they are the result of a self-consistency condition in the generative dynamics. We argue that the critical instability that arises from the phase transitions lies at the heart of their generative capabilities, which are characterized by a set of mean-field critical exponents. Finally, we show that the dynamic equation of the generative process can be interpreted as a stochastic adiabatic transformation that minimizes the free energy while keeping the system in thermal equilibrium.
@@ 1 Introduction
Generative modeling is a sub-field of machine learning concerned with the automatic generation of structured data such as images, videos and written language (Bond-Taylor et al.,, 2021). Generative diffusion models (Sohl-Dickstein et al.,, 2015), also known as score-based models, form a class of deep generative models that have demonstrated high performance in image (Ho et al.,, 2020; Song et al., 2021b, ), sound (Chen et al.,, 2020; Kong et al.,, 2020; Liu et al.,, 2023) and video generation (Ho et al.,, 2022; Singer et al.,, 2022). Diffusion models were first introduced in analogy with the physics of non-equilibrium statistical physics. The fundamental idea is to formalize generation as the probabilistic inverse of a forward stochastic process that gradually turns the target distribution $\phi(\boldsymbol{y})$ into a simple base distribution such as Gaussian white noise (Sohl-Dickstein et al.,, 2015; Song et al., 2021a, ). Recently, several works suggested that many of the dynamical properties of generative diffusion models can be understood using concepts such as spontaneous symmetry breaking (Raya and Ambrogioni,, 2023; Biroli and Mézard,, 2023; Biroli et al.,, 2024), and phase transitions (Biroli et al.,, 2024; Sclocchi et al.,, 2024). These theoretical and experimental results suggest a deep connection between generative diffusion and equilibrium phenomena.
In this paper, we outline a conceptual reformulation of generative diffusion models in the language of equilibrium statistical physics. We begin by defining a family of Boltzmann distributions over the noise-free states, which are interpreted as (unobservable) microstates during the diffusion process. In this picture, the Boltzmann weights are given by the conditional distributions of the noiseless data given the noisy state. We obtain a self-consistent equation of state for the system, which corresponds to the fixed-point equation of the generative dynamics. Moreover, we show that generative diffusion models can undergo second-order phase transitions of the mean-field type, corresponding the the generative spontaneous symmetry breaking phenomena fist discussed in (Raya and Ambrogioni,, 2023) and further studied in Biroli et al., (2024); Li and Chen, (2024) and in Sclocchi et al., (2024). Finally, we show that this mean-field theory can be seen as the thermodynamic limit of a multi-site system of coupled replicas. Based on this results, we derive a variant of the generative diffusion equations as the Brownian dynamics of a ’particle’ coupled on a large densely connected systems of replicated microstates, which offers a possible generalization of diffusion models beyond mean-field theory.
2 Contributions and Related Work
The main novel theoretical contributions of this paper are in the characterization of mean-field critical phase transitions in generative diffusion models, and its extension beyond mean-field theory. While the paper contains novel results, its aim is also pedagogical, as we wish to provide a self-consistent introduction for physicists to the study of generative diffusion. As such, we report known formulas and results from the literature, including the analysis scheme used in (Lucibello and Mézard,, 2024) and Biroli et al., (2024) for the analysis of memorization phenomena and the equivalence results for modern Hopfield networks given in (Ambrogioni,, 2023). Several of these formulas can also be found in recent work on stochastic localization (El A. et al.,, 2022; Huang et al.,, 2024), which has been shown to offer an elegant generalization of generative diffusion processes (Montanari,, 2023; Benton et al.,, 2024; Alaoui et al.,, 2023). In particular, the Boltzmann distributions given in Eq. 6 is equivalent to the tilted distributions given in (El A. et al.,, 2022).
@@ 3 Preliminaries on generative diffusion models
where the expectation is taken with respect to the target distribution $\phi(\boldsymbol{y})$. A generative model can then be obtained by "inverting" Eq. 1. The inverse equation is
@@ 3.1 Training diffusion models as denoising autoencoders
$\nabla\log{p_{t}(\boldsymbol{x})}\approx-(t\sigma^{2})^{-{1/2}}f^{*}(\boldsymbol{x}(t),t)~{}.$ (5)
@@ 4 Diffusion models as systems in equilibrium
The starting point for any model in statistical mechanics is the definition of the relevant microstates. In statistical physics, the microstate of a system is usually assumed to be an unobservable quantity. Given that we can observe a noise-corrupted data $\boldsymbol{x}^{*}(t)$, in a diffusion model the most obvious unobservable quantity of interest is the noise-free initial state $\boldsymbol{x}^{*}(0)=\boldsymbol{y}^{*}$. The next step is to define a Hamiltonian function on the set of microstates. We can do this by considering the conditional probability of the data $\boldsymbol{y}$ given a noisy state $\boldsymbol{x}_{t}$:
$p(\boldsymbol{y}\mid\boldsymbol{x},t)=\frac{1}{p_{t}(\boldsymbol{x})\sqrt{2\pi t\sigma^{2}}}e^{-\frac{{\left\lVert\boldsymbol{x}-\boldsymbol{y}\right\rVert}_{2}^{2}}{2t\sigma^{2}}+\log{\phi(\boldsymbol{y})}}=\frac{1}{Z(\boldsymbol{x},t)}e^{-\mathcal{H}(\boldsymbol{y};\boldsymbol{x},t)}~{},$ (6)
@@
The statistical properties of this ensemble determine the score function, which can be expressed as a Boltzmann average:
$\nabla\log{p_{t}(\boldsymbol{x})}=\beta(t)(\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}-\boldsymbol{x})~{},$ (9)
where $\beta(t)=(t\sigma^{2})^{-1}$ and
$\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}=\int\boldsymbol{y}~{}\frac{e^{-\mathcal{H}(\boldsymbol{y};\boldsymbol{x},t)}}{Z(\boldsymbol{x},t)}\text{d}\boldsymbol{y}~{}.$ (10)
Intuitively, this equation tells us that the score vector directs the system towards the posterior average $\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}$. As we shall see, studying the thermodynamics determined by these weights will allow us to understand several important qualitative and quantitative features of the generative dynamics. For example, as we will see in later sections, after a ’condensation’ phase transition the score will only depend on a small number of data-points, which can be detected by studying the concentration of the weights on a sub-exponential number of microstates.
The thermodynamic system defined by Eq. 6 does not have a true temperature parameter. However, the quantity $\sigma t$ plays a very similar role to temperature in classical statistical mechanics. Moreover, in the Hamiltonian given by Eq .39, the dynamic variable $\boldsymbol{x}$ is analogous to the external field term in magnetic systems, which can bias the distribution of microstates towards the patterns ’aligned’ in its direction. We can imagine $\boldsymbol{x}$ as being a "slower" thermodynamic variable that interacts (adiabatically) with the statistics of the microstates.
@@ 5 Free energy, magnetization and order parameters
$\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}=-\nabla\mathcal{F}(\boldsymbol{x},t)~{}.$ (20)
This formula suggests an analogy between diffusion models and magnetic systems in statistical physics. The noisy state $\boldsymbol{x}$ can be interpreted as an external magnetic field, which induces the state of "magnetization" $\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}$. In this analogy, a diffusion model is magnetized when its distribution is biased towards a sub-set of the possible microstates.
In physics, the ’external field’ variable $\boldsymbol{x}$ is usually assumed to be controlled by the experimenter. On the other hand, in generative diffusion models $\boldsymbol{x}$ is a dynamic variable that, under the reversed dynamics, is itself attracted towards $\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}$ by the drift term:
$\nabla\log{p_{t}(\boldsymbol{x})}=\beta(t)\left(\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}-\boldsymbol{x}\right)$ (21)
In other words, if we ignore the effect of the dispersion term, the state of the system is driven towards self-consistent points where $\boldsymbol{x}$ is equal to $\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}$. It is therefore interesting to study the self-consistency equation
@@
Readers familiar with statistical physics will recognize that Eq. 22 is formally identical to the self-consistency conditions used in the mean-field approximation, where the external field term in one location is assumed to be determined by the magnetization of all other locations. However, it is important to note that in the case of a diffusion model, this self-consistent coupling is not approximate, as it is a natural consequence of the dynamics. Nevertheless, the formal analogy implies that the thermodynamics of generative diffusion models is formally identical to the thermodynamics of mean-field models.
@@ 5.1 The susceptibility matrix
$\chi_{ij}(\boldsymbol{x},t)=\frac{\partial m_{i}}{\partial h_{j}}\bigg{\rvert}_{\boldsymbol{h}=0}~{},$ (24)
which tells us how sensitive the expected value is to changes in the noisy-state $\boldsymbol{x}(t)$. The susceptibility matrix is helpful in interpreting the dynamics of the generative denoising process as it informs us on how random fluctuations in each component of the state $\boldsymbol{x}(t)$ are propagated to the other components. For example, in the context of image generation, a random fluctuation of "green" at the bottom of an image can propagate to the rest, originating the image of a forest.
The susceptibility matrix can be re-expressed in temrms of connected correlation matrix (i.e. the covariance matrix) of the microstates under the Boltzmann distribution
$\chi_{ij}(\boldsymbol{x},t)=-\beta(t)\left(\langle\boldsymbol{y}\boldsymbol{y}^{T}\rangle_{\boldsymbol{x}}-\langle\boldsymbol{y}\rangle_{t,\boldsymbol{x}}\langle\boldsymbol{y}^{T}\rangle_{\boldsymbol{x}}\right)=-\beta(t)C_{ij}(\boldsymbol{x},t)$ (25)
We can now express the Jacobi matrix of the score function as follows
$J_{ij}(\boldsymbol{x})=\beta(t)\left(\chi_{ij}(\boldsymbol{x},t)-\delta_{j,k}\right)=-\beta(t)\delta_{j,k}-\beta(t)^{2}C_{ij}(\boldsymbol{x},t)~{}.$ (26)
@@ 6 Phase transitions and symmetry breaking
A spontaneous symmetry breaking happens when the trivial solution for the order parameter branches into multiple solutions at some critical time of $t_{c}$. This corresponds to the onset of multi-modality in the regularized free energy ~$\tilde{\mathcal{F}}(\boldsymbol{x},t)$, as visualized in Fig. 3 for a "four deltas" model. In thermodynamic systems, the symmetry breaking corresponds to a (second order) phase transition, which in this case can be detected by the divergence of several state variables around the critical point.
@@ 6.1 Generation and critical instability
7 Generation as as an adiabatic free energy descent process
where $\tilde{\mathcal{F}}$ is the free-energy plus a free potential term:
~$\tilde{\mathcal{F}}(\boldsymbol{x},t)=\mathcal{F}(\boldsymbol{x},t)+V(\boldsymbol{x})$ (32)
where $V(\boldsymbol{x})=\frac{1}{2}{\left\lVert\boldsymbol{x}\right\rVert}_{2}^{2}$. This can be seen as a form of adiabatic approximation, where the dynamics of the ’slow’ variable $\boldsymbol{x}$ is obtained by assuming that the system is maintained in thermal equilibrium along the diffusion trajectory. The symmetry breaking can now be detected as a change of shape in the regularized free energy, which transitions from a convex shape with a single global minimum to a more complex shape with potentially several meta-stable points (see Fig. 3). The reformulation of the dynamics in term of the gradient of a free energy allows us to interpret generative diffusion models as a kind of energy-based machine learning models (LeCun et al.,, 2006), as discussed in (Hoover et al.,, 2023) and (Ambrogioni,, 2023). The main difference is that the (free) energy is not learned directly but it is instead implicit in the learned score function. The connection suggests potential connection with the free energy principle in theoretical neuroscience, which is used to characterize the stochastic dynamics of biological neural systems (Friston,, 2010).
8 Beyond mean-field theory: A multi-site ’generative bath’ model
There results given in the previous sections suggest that generative diffusion models can be seen as a mean-field limit of a model with replicated microstates on $K$ ’sites’ coupled through long-range interactions. We denote the microstate in the $\mu$-th site as $\boldsymbol{y}^{\mu}$. Consider the following multi-site Hamiltonian:
$\mathcal{H}_{K}[\boldsymbol{y}^{1:K};\boldsymbol{h}]=\beta\left(\frac{w}{2}\sum_{\mu,\nu,~{}\mu\neq\nu}^{K}\boldsymbol{y}^{\mu}\cdot\boldsymbol{y}^{\nu}-\sum_{j}^{K}\boldsymbol{y}^{\mu}\cdot\boldsymbol{h}\right)-\sum_{j}\log{\psi(\boldsymbol{y}^{\mu})}$ (33)
where $\boldsymbol{y}^{1:K}$ denotes a multi-site configuration and $w$ is coupling weight parameter and $\boldsymbol{h}$ is an external field term. The re-normalized marginal energy $-\log{\psi(\boldsymbol{y}^{\mu})}$ is defined to fulfil the following constraint:
$\lim_{\beta\rightarrow\infty}\frac{1}{Z_{K}(\beta,\boldsymbol{h})}e^{-\beta\mathcal{H}_{K}[\boldsymbol{y}^{1:K};\boldsymbol{h}]}=\phi(\boldsymbol{y}^{1})\prod_{\mu=2}^{K}\delta(\boldsymbol{y}^{\mu}-\boldsymbol{y}^{1})~{}.$ (34)
In general, we have that $\log{\psi(\boldsymbol{y})}=\log{\phi(\boldsymbol{y})}$ when the possible microstates have a constant euclidean norm, while it will involve additional re-normalization terms in the general case. This condition ensures that the multi site coupling only affects the correlation between different sites (leading to perfect alignment for $\beta\rightarrow\infty$) while preserving the limiting marginal distributions. We refer to this thermodynamic system as a generative bath.
In the model, different replications of the microstates $\boldsymbol{y}_{j}$ (i.e. the noise-free data) exert mutual attractive couplings. Generation can be seen as a spontaneous symmetry breaking that the system undergoes when the temperature (i.e. the time) decreases, since at low temperatures all the microstates in all sites will align on the same pattern, resulting in a coherent observable average
$\bar{\boldsymbol{y}}=\frac{1}{K}\sum_{\mu}^{K}\boldsymbol{y}^{\mu}~{}.$ (35)
In the thermodynamic limit ($K\rightarrow 0$), the model converges to the self-consistent mean-field model discussed in the previous sections. This allows us to conceptualize the self-consistency condition implicit in the fixed-point equation of generative diffusion model as the result of an ideal multi-site coupling, which could result in new forms of neural implementations. This conceptualization opens the door for possible non-mean-field generalizations of generative diffusion characterized by short-range interactions, or disordered generalizations with random interactions. However, it is not clear if these extensions will have practical value.
Depending on the choice of the target distribution $\phi(\boldsymbol{y})$, the coupled model in Eq. 35 reduces to well known models in statistical mechanics such as the fully connected Ising model for the two-deltas distribution and the classical Heisenberg model for spherical distributions. However, the model becomes substantially more complex under more realistic distributions of the data.
8.1 Brownian dynamics in a ’generative bath’
The Hamiltonian defined in Eq. 35 specifies an equilibrium system that, in the thermodynamic limit, shares the same self-consistent criticality of the fixed-point equation of generative diffusion models. In this section, we will derive from first principles a generative stochastic dynamics similar to the generative equation given in Eq. 3. The idea is to consider a ’Brownian particle’ $\boldsymbol{x}(t)$ coupled to the multi-site system of microstates. We define the random force as
$F(\boldsymbol{x},t)=\frac{1}{H}\sum_{\mu}^{H}\left(\boldsymbol{y}^{\mu}-\boldsymbol{x}\right)~{},$ (36)
where $H$ is the number of sites that are coupled to $\boldsymbol{x}$. In contrast to the distribution in Eq. 6, we assume that $\boldsymbol{x}$ does not exert any effect on the equilibrium systems itself, which instead undergoes symmetry breaking events due to its own internal coupling between sites. In other word, in this formulation the state $\boldsymbol{x}(t)$ is now passively controlled by the statistical fluctuations in the equilibrium system.
If we assume that the force in Eq. 36 is applied at each infinitesimal time interval and that its time-scale is much faster that the motion of $\boldsymbol{x}(t)$, the Brownian dynamics follows the following (reversed) Langevin equation
$\boldsymbol{x}(t-\text{d}t)-\boldsymbol{x}(t)=-\left(\langle\bar{\boldsymbol{y}}\rangle_{t,\boldsymbol{x}\rightarrow 0}-\boldsymbol{x}(t)\right)\frac{\text{d}t}{t}+\frac{1}{\sqrt{H}}B(\boldsymbol{x},t)\boldsymbol{w}(t)\sqrt{\frac{\text{d}t}{t}}~{},$ (37)
where $B(\boldsymbol{x},t)=C^{1/2}(\boldsymbol{x})$ is a matrix square root of the pure state covariance matrix:
$C(\boldsymbol{x})=\left(\langle\bar{\boldsymbol{y}}\bar{\boldsymbol{y}}^{T}\rangle_{t,\boldsymbol{x}\rightarrow 0}-\langle\bar{\boldsymbol{y}}\rangle_{t,\boldsymbol{x}\rightarrow 0}\langle\bar{\boldsymbol{y}}^{T}\rangle_{t,\boldsymbol{x}\rightarrow 0}\right)~{}.$ (38)
The $1/t$ scaling in the differential are introduced to have the reversed diffusion ends at the finite time $t=0$, which is equivalent to a logarithmic change of coordinate in the time variable.
The Boltzmann expectation is taken with respect to the multi-site ensemble given in Eq. 35 in the limit of a vanishing external field alligned to $\boldsymbol{x}$. This is done in order to isolate the appropriate ’pure state’ from the Boltzmann average, since after a spontaneous symmetry breaking only one branch of the distribution should affect the particle. In fact, after a symmetry breaking phase transition, the Boltzmann distribution splits into two or more modes corresponding to the possible states with broken symmetry.
8.2 The two delta model revisited
In the "two deltas" model, the Hamiltonian of the generative bath" is just
$\mathcal{H}_{K}[y^{1:K};h]=\beta\left(\frac{w}{2}\sum_{\mu,\nu,~{}\mu\neq\nu}^{K}y^{\mu}y^{\nu}-h\sum_{j}^{K}y^{\mu}\cdot\right)$ (39)
with the restriction that that $y^{\mu}\in\{\pm 1\}$. This is simply the Hamiltonian of a fully connected Ising model with uniform coupling weights. In the thermodynamic limit $K\rightarrow\infty$, the model therefore reduces to the mean-field Curie-Weiss model that we have already discussed. In this case, the pure state magnetisation are the stable solutions $m_{t}$ of the self-consistency equation $m=\tanh{(m+h)/(t\sigma^{2})}$, which is identically equal to zero for $t\sigma^{2}>1$ and it has two branches that cannot be expressed in closed-form in the low-temperature regime. The instantaneous variance of the Brownian generative dynamics is given by $T\partial m_{t}/\partial h$, which is equal to $1/(1-(t\sigma^{2})^{-1})$ in the high temperature phase and to $T(1-m_{t}^{2})/m_{t}^{2}$ in the low temperature phase. Note that the variance diverges at $t\sigma^{2}=1$ due to the critical phase transition and that it vanishes for $t\rightarrow 0$ as the system fully aligns in one of the two possible pure states.